# MetaWFN: A Platform for Unified Implementation of Many-Electron Wavefunctions

### v1

@@ 5 $e_{ij,kj}$ $k>i>j$ d4
Figure 5: Graphical representations of Eq. (10) (a) and Eq. (14) (b).
@@ 2.2 Unified Handling of MPBFs
To achieve a unified handling of MPBFs, the CFGs are taken as the organization unit in MetaWFN. An CFG $|I\rangle$, defined by a set of occupation numbers $\in\{0,1,2\}$ for the $n$ spatial orbitals in the spin-free case or $n$ Kramers pairs of spinors in the spin-dependent case, can be represented by a binary string composed of an array of 64-bit unsigned integers (2 bits for each spatial orbit/Kramers pair), viz.
@@
where $n_{p}^{q}=0,1,2$ are represented by $(00)_{2},(01)_{2}$, and $(11)_{2}$, respectively. Here, the number of bit 1 coincides with the occupation number, thereby enabling efficient bitwise operations on CFGs13, 14. An CFG $|I\rangle$ can generate a number of CSFs/DETs denoted as $|I\mu\rangle$, where $\mu$ distinguishes different spin-coupling schemes for CSFs or distributions of open-shell spins/Kramers partners for DETs.
@@ 2.3 Unified Construction of Hamiltonian Matrix
The HMEs $\langle I\mu|H|J\nu\rangle$ involve contractions between the integrals and BCCs, the reuse of which must be maximized in sCI methods due to the lack of structures. To facilitate this, we classify the HMEs into three ‘matrix element categories’ (MEC) corresponding to zero-, single-, and two-electron differences between CFGs $I$ and $J$, and meanwhile classify the two-electron integrals into seven ‘integral categories’ (IntC$\in[1,7]$ in Table 1). Each MEC/IntC involves only a subset of the integrals determined by the CFG pair $(I,J)$ rather than by the individual MPBFs. For the cases with zero- and one-electron differences, the HMEs can be simplified by introducing a closed-shell reference CFG ${\omega}$ in both the spin-free13 and spin-dependent7 cases.
As for the reuse of the BCCs, a key observation lies in that doubly occupied or unoccupied orbitals common to the CFG pair $(I,J)$ can be permuted freely with other orbitals without changing the BCCs, provided that the Yamaguchi–Kotani phase is employed in the case of CSFs49. Consequently, they can be deleted, leading to a reduced CFG pair $(\tilde{I},\tilde{J})$ spanned by the reduced set of orbitals $\tilde{r}\in[\tilde{1},\tilde{N}]$, with $\tilde{N}$ being the number of common singly occupied orbitals and those orbitals with different occupations in $\tilde{I}$ and $\tilde{J}$. In this way, the BCCs $\langle I\mu|E_{ij}|J\nu\rangle$ and $\langle I\mu|e_{ij,kl}|J\nu\rangle$ can simply be calculated as $\langle\tilde{I}\mu|E_{\tilde{i}\tilde{j}}|\tilde{J}\nu\rangle$ and $\langle\tilde{I}\mu|e_{\tilde{i}\tilde{j},\tilde{k}\tilde{l}}|\tilde{J}\nu\rangle$, respectively. The orbital sequence and occupation pattern of $(\tilde{I},\tilde{J})$ can be encoded into a reduced occupation table (ROT)13 as the internal representation in MetaWFN. Different CFG pairs but of the same ROT share the same BCCs. Moreover, due to Hermiticity of the BCCs, only those ROTs with $\tilde{I}\geq\tilde{J}$ are needed in practical applications, thereby achieving optimal reuse of the BCCs.
@@
    Classify the given set of CFG pairs $\{(I,J)\}$ according to the MECs and then establish the ROTs for each MEC;
@@
    Sort CFG pairs $\{(I,J)\}$ according to the ROTs;
@@
    Evaluate the BCCs on the fly across all CFG pairs of the same ROT, and then contract with different integrals for different HMEs.
In summary, different Hamiltonians, different MPBFs, and unstructured HMEs can be treated in unified manners within MetaWFN.
@@ 2.4 Selection
The aim of selection is to find iteratively an improved variational space for the expansion of the wavefunction. It involes four key components, i.e., ranking, diagonalization (of $P_{1}^{[i]}$ in Fig. 1), pruning, and termination. To simplify the notation, the variational space and its FOIS are denoted here as $P$ and $Q$, respectively. The former gives rise to the target states $\{E^{(0)}_{k},|\Psi^{(0)}_{k}\rangle=\sum_{J\nu\in P}|J\nu\rangle C_{\nu,k}^{J}\}$.
@@
$f\left(|I\mu\rangle\right)=\max_{k}\left|\frac{\sum_{|J\nu\rangle P}H_{\mu\nu}^{IJ}C^{J}_{\nu,k}}{E^{(0)}_{k}-H_{\mu\mu}^{II}}\right|,\quad H_{\mu\nu}^{IJ}=\langle I\mu|H|J\nu\rangle$ (17)
@@
Note in passing that the perturbative estimate (18) of $\epsilon\left(|I\mu\rangle\right)$ can also be replaced57 with that obtained by diagonalizing the Hamiltonian over the two-state system $\{\Psi_{0},|I\mu\rangle\}$. Such ‘CIPSI criterion’ can be used in different ways when constructing the $P$ space iteratively. For instance, all $\{|I\mu\rangle\}\in Q$ with $f\left(|I\mu\rangle\right)$ larger than a preset threshold $C_{\mathrm{Q}}$ are retained in the variational space $P$ at each iteration of CIPSI (configuration interaction with perturbative selection made iteratively)58. It is obvious that the final size of $P$ (denoted as $|P|$) is determined by the threshold $C_{\mathrm{Q}}$ itself. A variant59 of CIPSI is that those $\{|I\mu\rangle\}\in Q$ with the largest $\epsilon\left(|I\mu\rangle\right)$ are put into $P$ until the target size $|P|$ has reached. An additional constraint is invoked in ASCI (adaptive sampling configuration interaction)27, where except for $|P|$, $|Q|$ is also kept fixed throughout the iterative selections, by making use of $f\left(|I\mu\rangle\right)$. A different strategy is adopted in ACI (adaptive configuration interaction)24, where those $\{|I\mu\rangle\}\in Q$ of the smallest $\epsilon\left(|I\mu\rangle\right)$ are excluded as long as the sum of their $\epsilon\left(|I\mu\rangle\right)$ is smaller than the target accuracy of energy. At variance with these implementations which feature two spaces ($P$ and $Q$), ICE (iterative configuration expansion)32 adopts the three-space concept60, where the $P$ space is split into two subspaces spanned by the generators and the rest (spectators), which have weights larger and smaller than a threshold, respectively. Only the former are allowed to access perturbers $\{|I\mu\rangle\}\in Q$, which are included in the $P$ space only when their $\epsilon\left(|I\mu\rangle\right)$ are larger than a second threshold.
Noticing that the CIPSI criterion (17)/(18) is computationally too expensive (due to the summation in the numerator), a drastically simplified criterion was introduced in SHCI (semi-stochastic heat-bath configuration interaction)16, viz.
@@
It follows that those $|I\mu\rangle\in Q$ with $|H^{IJ}_{\mu\nu}|<\eta_{Q}/\max_{k}|C^{J}_{\nu,k}|$ are never touched (cf. Fig. 6), provided that the integrals are pre-sorted in descending order of their magnitudes. However, such integral-driven selection usually leads to a variational space $P$ that is much less compact14 than that by using the coefficient/energy-driven selection (17)/(18). To remedy this but retaining the efficiency, the good of the integral- and coefficient-driven selections is combined in iCIPT213, 14, by means of the following boolean function $f\left(|I\mu\rangle,|J\rangle,C_{\min}\right)$:
@@
Literally, for case (A), loop over $|I\mu\rangle$ in $Q$ and evaluate $H_{\mu\nu}^{IJ}$ for all CSFs $|J\nu\rangle\in P$. If $\max_{\nu,k}\left(\left|H_{\mu\nu}^{IJ}C^{J}_{\nu,k}\right|\right)$ is larger than $C_{\min}$ then evaluate $H_{\mu\mu}^{II}$; otherwise discard $|I\mu\rangle$. If $\max_{\nu,k}\left(\left|\frac{H_{\mu\nu}^{IJ}C^{J}_{\nu,k}}{E^{(0)}_{k}-H_{\mu\mu}^{II}}\right|\right)$ is larger than $C_{\min}$ then $|I\mu\rangle$ is selected. As for case (B), only those doubly excited CFGs $|I\rangle$ with estimated upper bounds $\tilde{H}^{IJ}$ (cf. Table 1) larger than $C_{\min}/\max_{\nu,k}$ $|C^{J}_{\nu,k}|$ need to be generated (i.e., those unimportant ones are never touched, see Fig. 6). For such $\{|I\rangle\}$, the remaining step is the same as case (A).
It should be noted that not all CSFs/DETs have appreciable weights in the wavefunction ($\Psi_{1}^{[i]}$ in Fig. 1), even with the above ranking. Therefore, those CFS/DETs of very small coefficients should be deleted to end up with a compact varitional space ($P_{2}^{[i]}$ in Fig. 1) that is affordable for subsequent PT2 correction. Such a pruning step is invoked in ASCI (by retaining only a fixed number of DETs of the largest coefficients)27, ACI (by enforcing that the sum of the squared coefficients of DETs is smaller than a preset threshold)24, and iCIPT2 (by retaining all CSFs with coefficients larger than $C_{\mathrm{min}}$ in absolute value)13, but not in CIPSI58, 59, SHCI16, and ICE32. While this is not very serious for CIPSI and ICE, it is not the case for SHCI due to the use of the very loose criterion (19). It has been shown14 that typically 90% of the DETs selected according to the criterion (19) can be pruned away, meaning that a substantial time is wasted in the diagonalization. Speaking of the overall efficiency, the whole selection procedure with the iCI criterion (20)/(21) is about five times cheaper, whereas that with the CIPSI (17) or SHCI (19) criterion is about two times more expensive than the PT2 correction step (for more details, see Ref. 14).
@@
Finally, some words on the diagonalization are necessary. As well known, the best way to make full use of the sparsity of the Hamiltonian matrix $\mathbf{H}$ is through the matrix-vector products, $\bm{\sigma}_{i}=\mathbf{H}\mathbf{X}_{i}$, which require only nonzero elements of $\mathbf{H}$ (which are never stored but are recalculated whenever needed). The crucial point here is how to choose the trial (column) vectors $\{\mathbf{X}_{i}\}_{i=1}^{N_{b}}$, so as to expand the target roots accurately even if the space $\mathbf{X}=\mathrm{span}\{\mathbf{X}_{i}\}$ is very incomplete. Unfortunately, most schemes work with a search space $\mathbf{X}$ whose dimension $N_{b}$ increases along the iterations, thereby leading to large memory consumptions. To avoid this, iCIPT2 adopts the iterative vector interaction (iVI) approach61, 62, which has the following merits: (1) it adopts well defined trial vectors (by perturbation theory, intermediate Hamiltonian, conjugate gradient, etc.); (2) it works with a fixed-dimensional search space, with the dimension determined automatically by the number of target roots; (3) it can access directly interior roots without touching none of the lower roots.
@@ 2.5 Constraint-Based ENPT2
As long as the iteratively determined variational space $P$ is good enough, the remaining dynamic correlation can be accounted for accurately at the lowest-order perturbation theory. To be more general, the three-space concept60, 32 can be adopted here. That is, the $P$ space is split into a primary ($P_{m}$) and a secondary ($P_{s}$) subspace, and only the former is allowed to connect the external space $Q=1-P$. The eigenpairs $\{\tilde{E}^{(0)}_{k},|\tilde{\Psi}^{(0)}_{k}\rangle=\sum_{J\nu\in P_{m}}|J\nu\rangle C_{\nu,k}^{J}\}$ of the $P_{m}$ space are to be corrected by ENPT2 in a state-specific manner, viz.
@@
    How to establish the connections between the CFGs $I\in Q$ with those interacting CFGs $J\in P_{m}$?
@@
To address these issues, note first that the single and double excitations from the CSFs/DETs of the $P_{m}$ space in general span a space $W$ that is larger than $Q$, for some of them may belong to $P_{m}$ or $P_{s}$. That is, the net $Q$ space can only be obtained by removing the $P$-space CSFs/DETs from $W$. However, this requires a very slow double-check process. Fortunately, this issue can be bypassed by rewriting Eq. (23) as29
@@
Inspired by the idea of constraint PT229, we have developed a pipelined, residue-based constraint PT2 algorithm (see Fig. 7; for more details see Ref. 14):
@@
    Split the $W$ space into disjoint subspaces $\{W_{i}\}$ (by constraints consisting of $L_{c}$ highest occupied orbitals and corresponding occupation numbers) and dynamically distribute $\{W_{i}\}$ across threads or processes;
@@
    Determine the memory to store the CFG connection information for each $W_{i}$. If $W_{i}$ is too large, split it into smaller subspaces until the memory requirement is met;
@@
    Generate the CFG connection information for each $W_{i}$. If the consumed memory does not reach the threshold, pick a new subspace and return to Step 2.
@@
    When the memory for storing the CFG connection information reaches the given threshold, sort the connection information to determine the CFGs $I$ in $W_{i}$ and those interacting CFGs $J$ in $P$;
@@
    For each CFG $I$ in $W_{i}$, allocate memory to store the denominator and numerator for the PT2 correction and generate the corresponding ROT for each connected CFG pair $(I,J)$. If the memory at this step reaches a predefined threshold, then launch the calculation to get the PT2 correction from all involved CFGs. Here, the BCCs and molecular integrals can be used repeatedly for optimal efficiency.
@@
    After finishing $\tilde{E}_{k}^{(2)}[W]$, repeat the same for $\tilde{E}_{k}^{(2)}[P]$.
@@
Figure 7: Workflow of pipelined, residue-based constraint ENPT2 in MetaWFN. Green boxes represent the start and end points of the algorithm. Blue boxes indicate steps that must be executed either in single-thread mode or multi-thread mode with locks. The yellow box serves as a pool for storing connection records and managing memory usage across threads. White boxes denote steps that can be performed independently by each thread, without the need for synchronization.

### v2

@@ 5 $e_{ij,kj}$ $k>i>j$ d4
$i$$i$ $i$$i$ $i$$i$$j$$j$ $i$$i$$j$$j$
(a) w1 (b) w2 (c) w3 (d) b7=c7
Figure 2: Diagrammatic representation of $H_{1}^{0}$ (a) and $H_{2}^{0}$ (b-d). The naming of the diagrams follows Ref. 58
$i$$j$ $j$$i$ $i$$i$$j$ $j$$i$$i$ $i$$j$$j$ $j$$i$$j$
(a) s1 (b) s2 (c) s3 (d) s4 (e) s5 (f) s6
$k$$j$$i$$k$ $k$$j$$i$$k$ $i$$k$$k$$j$ $i$$k$$k$$j$ $i$$k$$k$$j$ $i$$k$$k$$j$
(g) s7 (h) s8 (i) s9 (j) s10 (k) s11 (l) s12
$k$$k$$i$$j$ $k$$k$$i$$j$ $k$$k$$i$$j$ $k$$k$$i$$j$ $k$$k$$i$$j$ $k$$k$$i$$j$
(m) b6 (n) c6 (o) b4 (p) c4 (q) a2 (r) d2
Figure 3: Diagrammatic representation of $H_{1}^{1}$ (a-b) and $H_{2}^{1}$ (c-r). The naming of the diagrams follows Ref. 58
$i$$k$$j$$l$ $i$$k$$j$$l$ $j$$l$$i$ $i$$k$$j$$l$ $i$$j$$k$ $j$$i$
(a) a1 (b) a3 (c) a4 (d) a5 (e) a6 (f) a7
$i$$k$$j$$l$ $j$$l$$i$ $i$$k$$j$$l$ $i$$k$$j$$l$
(g) b1 (h) b2 (i) b3 (j) b5
$i$$k$$j$$l$ $i$$k$$j$ $i$$k$$j$$l$ $i$$k$$j$$l$
(k) c1 (l) c2 (m) c3 (n) c5
$i$$k$$j$$l$ $i$$k$$j$$l$ $i$$k$$j$ $i$$k$$j$$l$ $j$$l$$i$ $j$$i$
(o) d1 (p) d3 (q) d4 (r) d5 (s) d6 (t) d7
Figure 4: Diagrammatic representation of $H_{2}^{2}$. The naming of the diagrams follows Ref. 58
Figure 5: Graphical representations of (a) product of segment values $W_{r}$ ($r\in[1,n]$) in Eq. (10) and (b) product of transfer matrices $\mathbf{E}(i)$ ($i\in[1,n]$) in Eq. (14).
@@ 2.2 Unified Handling of MPBFs
To achieve a unified handling of MPBFs, the oCFGs are taken as the organization unit in MetaWave. An oCFG $|I\rangle$, defined by a set of occupation numbers $\in\{0,1,2\}$ for the $n$ spatial orbitals in the spin-free case or $n$ Kramers pairs of spinors in the spin-dependent case, can be represented by a binary string composed of an array of 64-bit unsigned integers (2 bits for each spatial orbit/Kramers pair), viz.
@@
where $n_{p}^{q}=0,1,2$ are represented by $(00)_{2},(01)_{2}$, and $(11)_{2}$, respectively. Here, the number of bit 1 coincides with the occupation number, thereby enabling efficient bitwise operations on oCFGs16, 17. An oCFG $|I\rangle$ can generate a number of CSFs/DETs denoted as $|I\mu\rangle$, where $\mu$ distinguishes different spin-coupling schemes for CSFs or distributions of open-shell spins/Kramers partners for DETs.
@@ 2.3 Unified Construction of Hamiltonian Matrix
The HMEs $\langle I\mu|H|J\nu\rangle$ involve contractions between the integrals and BCCs, the reuse of which must be maximized in sCI methods due to the lack of structures (natural connections). To facilitate this, we classify the HMEs into three ‘matrix element categories’ (MEC) corresponding to zero-, single-, and two-electron differences between oCFGs $|I\rangle$ and $|J\rangle$, and meanwhile classify the two-electron integrals into seven ‘integral categories’ (IntC$\in[1,7]$ in Table 1). Each MEC/IntC involves only a subset of the integrals determined by the oCFG pair $(I,J)$ rather than by the individual MPBFs. For the cases with zero- and one-electron differences, the HMEs can be simplified by introducing a closed-shell reference oCFG $|\omega\rangle$ in both the spin-free16 and spin-dependent7 cases.
As for the reuse of the BCCs, a key observation lies in that doubly occupied or unoccupied orbitals common to the oCFG pair $(I,J)$ can be permuted freely with other orbitals without changing the BCCs, provided that the Yamaguchi–Kotani phase is employed in the case of CSFs58. Consequently, they can be deleted, leading to a reduced oCFG pair $(\tilde{I},\tilde{J})$ spanned by the reduced set of orbitals $\tilde{r}\in[\tilde{1},\tilde{N}]$, with $\tilde{N}$ being the number of common singly occupied orbitals and those orbitals with different occupations in $|\tilde{I}\rangle$ and $|\tilde{J}\rangle$. In this way, the BCCs $\langle I\mu|E_{ij}|J\nu\rangle$ and $\langle I\mu|e_{ij,kl}|J\nu\rangle$ can simply be calculated as $\langle\tilde{I}\mu|E_{\tilde{i}\tilde{j}}|\tilde{J}\nu\rangle$ and $\langle\tilde{I}\mu|e_{\tilde{i}\tilde{j},\tilde{k}\tilde{l}}|\tilde{J}\nu\rangle$, respectively. The orbital sequence and occupation pattern of $(\tilde{I},\tilde{J})$ can be encoded into a reduced occupation table (ROT)16 as the internal representation in MetaWave. Different oCFG pairs but of the same ROT share the same BCCs. Moreover, due to Hermiticity of the BCCs, only those ROTs with $\tilde{I}\geq\tilde{J}$ are needed in practical applications, thereby achieving optimal reuse of the BCCs.
@@
    Classify the given set of oCFG pairs $\{(I,J)\}$ according to the MECs and then establish the ROTs for each MEC;
@@
    Sort oCFG pairs $\{(I,J)\}$ according to the ROTs;
@@
    Evaluate the BCCs on the fly across all oCFG pairs sharing the same ROT, and then contract with different integrals for different HMEs.
In summary, different Hamiltonians, different MPBFs, and unstructured HMEs can be treated in unified manners within MetaWave.
@@ 2.4 Selection
The aim of selection is to find iteratively an improved variational space for the expansion of the wavefunction. It involes four key components, i.e., ranking, diagonalization (of $P_{1}^{[i]}$ in Fig. 1), pruning, and termination. To simplify the notation, the variational space and its first-order interacting space (FOIS) are denoted here as $P$ and $Q$, respectively. The former gives rise to the target states $\{E^{(0)}_{k},|\Psi^{(0)}_{k}\rangle=\sum_{J\nu\in P}|J\nu\rangle C_{\nu,k}^{J}\}$.
@@
$f\left(|I\mu\rangle\right)=\max_{k}\left|\frac{\sum_{|J\nu\rangle\in P}H_{\mu\nu}^{IJ}C^{J}_{\nu,k}}{E^{(0)}_{k}-H_{\mu\mu}^{II}}\right|,\quad H_{\mu\nu}^{IJ}=\langle I\mu|H|J\nu\rangle$ (17)
@@
Note in passing that the perturbative estimate (18) of $\epsilon\left(|I\mu\rangle\right)$ can also be replaced66 with that obtained by diagonalizing the Hamiltonian over the two-state system $\{\Psi_{0},|I\mu\rangle\}$. Such ‘CIPSI criterion’ can be used in different ways when constructing the $P$ space iteratively. For instance, all $\{|I\mu\rangle\}\in Q$ with $f\left(|I\mu\rangle\right)$ larger than a preset threshold $C_{\mathrm{Q}}$ are retained in the variational space $P$ at each iteration of CIPSI (configuration interaction with perturbative selection made iteratively)67. It is obvious that the final size of $P$ (denoted as $|P|$) is determined by the threshold $C_{\mathrm{Q}}$ itself. A variant68 of CIPSI is that those $\{|I\mu\rangle\}\in Q$ with the largest $\epsilon\left(|I\mu\rangle\right)$ are put into $P$ until the target size $|P|$ has reached. An additional constraint is invoked in ASCI (adaptive sampling configuration interaction)30, where except for $|P|$, $|Q|$ is also kept fixed throughout the iterative selections, by making use of $f\left(|I\mu\rangle\right)$. A different strategy is adopted in ACI (adaptive configuration interaction)27, where those $\{|I\mu\rangle\}\in Q$ of the smallest $\epsilon\left(|I\mu\rangle\right)$ are excluded as long as the sum of their $\epsilon\left(|I\mu\rangle\right)$ is smaller than the target accuracy of energy. At variance with these implementations which feature two spaces ($P$ and $Q$), ICE (iterative configuration expansion)35 adopts the three-space concept69, where the $P$ space is split into two subspaces spanned by the generators and the rest (spectators), which have weights larger and smaller than a threshold, respectively. Only the former are allowed to access perturbers $\{|I\mu\rangle\}\in Q$, which are included in the $P$ space only when their $\epsilon\left(|I\mu\rangle\right)$ are larger than a second threshold.
Noticing that the CIPSI criterion (17)/(18) is computationally too expensive (due to the summation in the numerator), a drastically simplified criterion was introduced in SHCI (semi-stochastic heat-bath configuration interaction)19, viz.
@@
It follows that those $|I\mu\rangle\in Q$ with $|H^{IJ}_{\mu\nu}|<\eta_{Q}/\max_{k}|C^{J}_{\nu,k}|$ are never touched (cf. Fig. 6), provided that the integrals are pre-sorted in descending order of their magnitudes. However, such integral-driven selection usually leads to a variational space $P$ that is much less compact17 than that by using the coefficient/energy-driven selection (17)/(18). To remedy this but retaining the efficiency, the good of the integral- and coefficient-driven selections is combined in iCIPT216, 17, by means of the following boolean function $f\left(|I\mu\rangle,|J\rangle,C_{\min}\right)$:
@@
Literally, for case (A), loop over $|I\mu\rangle$ in $Q$ and evaluate $H_{\mu\nu}^{IJ}$ for all CSFs $|J\nu\rangle\in P$. If $\max_{\nu,k}\left(\left|H_{\mu\nu}^{IJ}C^{J}_{\nu,k}\right|\right)$ is larger than $C_{\min}$ then evaluate $H_{\mu\mu}^{II}$; otherwise discard $|I\mu\rangle$. If $\max_{\nu,k}\left(\left|\frac{H_{\mu\nu}^{IJ}C^{J}_{\nu,k}}{E^{(0)}_{k}-H_{\mu\mu}^{II}}\right|\right)$ is larger than $C_{\min}$ then $|I\mu\rangle$ is selected. As for case (B), only those doubly excited oCFGs $|I\rangle$ with estimated upper bounds $\tilde{H}^{IJ}$ (cf. Table 1) larger than $C_{\min}/\max_{\nu,k}$ $|C^{J}_{\nu,k}|$ need to be generated (i.e., those unimportant ones are never touched, see Fig. 6). For such $\{|I\rangle\}$, the remaining step is the same as case (A).
It should be noted that not all CSFs/DETs have appreciable weights in the wavefunction ($\Psi_{1}^{[i]}$ in Fig. 1), even with the above ranking. Therefore, those CSF/DETs of very small coefficients should be deleted to end up with a compact varitional space ($P_{2}^{[i]}$ in Fig. 1) that is affordable for subsequent PT2 correction. Such a pruning step is invoked in ASCI (by retaining only a fixed number of DETs of the largest coefficients)30, ACI (by enforcing that the sum of the squared coefficients of DETs is smaller than a preset threshold)27, and iCIPT2 (by retaining all CSFs with coefficients larger than $C_{\mathrm{min}}$ in absolute value)16, but not in CIPSI67, 68, SHCI19, and ICE35. While this is not very serious for CIPSI and ICE, it is not the case for SHCI due to the use of the very loose criterion (19). It has been shown17 that typically 90% of the DETs selected according to the criterion (19) can be pruned away, meaning that a substantial time is wasted in the diagonalization. Speaking of the overall efficiency, the whole selection procedure with the iCI criterion (20)/(21) is about five times cheaper, whereas that with the CIPSI (17) or SHCI (19) criterion is about two times more expensive than the PT2 correction step (for more details, see Ref. 17).
@@
Finally, some words on the diagonalization are necessary. As well known, the best way to make full use of the sparsity of the Hamiltonian matrix $\mathbf{H}$ is through the matrix-vector products, $\bm{\sigma}_{i}=\mathbf{H}\mathbf{X}_{i}$, which require only nonzero elements of $\mathbf{H}$ (which are never stored but are recalculated whenever needed). The crucial point here is how to choose the trial (column) vectors $\{\mathbf{X}_{i}\}_{i=1}^{N_{b}}$, so as to expand the target roots accurately even if the space $\mathbf{X}=\mathrm{span}\{\mathbf{X}_{i}\}$ is very incomplete. Unfortunately, most schemes work with a search space $\mathbf{X}$ whose dimension $N_{b}$ increases along the iterations, thereby leading to large memory consumptions. To avoid this, iCIPT2 adopts the iterative vector interaction (iVI) approach70, 71, which has the following merits: (1) it adopts well defined trial vectors (by perturbation theory, intermediate Hamiltonian, conjugate gradient, etc.); (2) it works with a fixed-dimensional search space, with the dimension determined automatically by the number of target roots; (3) it can access directly interior roots without touching none of the lower roots.
@@ 2.5 Constraint-Based ENPT2
As long as the iteratively determined variational space $P$ is good enough, the remaining dynamic correlation can be accounted for accurately at the lowest-order perturbation theory. To be more general, the three-space concept69, 35 can be adopted here. That is, the $P$ space is split into a primary ($P_{m}$) and a secondary ($P_{s}$) subspace, and only the former is allowed to connect the external space $Q=1-P$. The eigenpairs $\{\tilde{E}^{(0)}_{k},|\tilde{\Psi}^{(0)}_{k}\rangle=\sum_{J\nu\in P_{m}}|J\nu\rangle C_{\nu,k}^{J}\}$ of the $P_{m}$ space are to be corrected by ENPT2 in a state-specific manner, viz.
@@
    How to establish the connections between the oCFGs $|I\rangle\in Q$ and those interacting oCFGs $|J\rangle\in P_{m}$?
@@
To address these issues, note first that the single and double excitations from the CSFs/DETs of the $P_{m}$ space in general span a space $W$ that is larger than $Q$, for some of them may belong to $P_{m}$ or $P_{s}$. That is, the net $Q$ space can only be obtained by removing the $P$-space CSFs/DETs from $W$. However, this requires a very slow double-check process. Fortunately, this issue can be bypassed by rewriting Eq. (23) as32
@@
Inspired by the idea of constraint-based PT232, we have developed a pipelined, constraint-based PT2 algorithm (see Fig. 7; for more details see Ref. 17):
@@
    Split the $W$ space into disjoint subspaces $\{W_{i}\}$ (by constraints consisting of $L_{c}$ highest occupied orbitals and corresponding occupation numbers) and dynamically distribute $\{W_{i}\}$ across threads or processes. Information for generating $W_{i}$ is stored in a pool that is shared by all threads.
@@
    Picks up a subspace $W_{i}$ and determine the memory to store the oCFG connection information. If $W_{i}$ is too large, split it into smaller subspaces until the memory requirement is met and push those smaller subspaces back to the pool and return to Step 2.
@@
    Generate the oCFG connection information for each $W_{i}$. If the consumed memory does not reach the threshold, pick a new subspace and return to Step 2.
@@
    Sort the connection information to determine the oCFGs $|I\rangle$ in $W_{i}$ and those interacting oCFGs $|J\rangle$ in $P$. Here, those doubly excited oCFGs $|I\rangle$ with estimated upper bounds $\tilde{H}^{IJ}$ (cf. Table 1) smaller than $Q_{\min}/\max_{\nu,k}$ $|C^{J}_{\nu,k}|$ will be deleted, so as to enhance the efficiency.
@@
    For each oCFG $|I\rangle$ in $W_{i}$, allocate memory to store the numerator and denominator in Eq. (25) and generate the corresponding ROT for each connected oCFG pair $(I,J)$. If the memory at this step reaches a predefined threshold, then launch the calculation to get the PT2 correction from all involved oCFGs. Here, use of the sparsity of $\mathbf{H}^{IJ}$ as well as repeated use of the BCCs and molecular integrals should be made for optimal efficiency. If the memory does not exceed the predefined threshold, then pick up a new subspace and go back to Step 2.
@@
    After finishing $\tilde{E}_{k}^{(2)}[W]$ in Eq. (25), calculate $\tilde{E}_{k}^{(2)}[P]$ in Eq. (26)/(28) and $\tilde{E}_{k}^{(2)}[Q]$ in Eq. (24).
@@
Figure 7: Workflow of pipelined, residue-based constraint-based ENPT2 in MetaWave. Green boxes represent the start and end points of the algorithm. Pool of subspaces store the those $W_{i}$ that to be generated, which is shared for each thread. Yellow boxes serves as a pool for storing connection or calculation information, which is private for each thread. Connection information refers to connected oCFG pairs and caluclation information refers to the ROT for each connected oCFG pair. Blue boxes indicate steps that must be executed either in single-thread mode or multi-thread mode with locks. White boxes denote steps that can be performed independently by each thread.